OpenStudy (anonymous):

The sum of two numbers is 52. The greater number is 4 more than the smaller number. Which equation can be used to solve for the smaller number? x − (x + 4) = 52 x + (x + 4) = 52 x(x + 4) = 52 x(x − 4) = 52

OpenStudy (anonymous):

I need help with this now, or I will fail this assignment

OpenStudy (zaynab123):

52-(x+4)

OpenStudy (anonymous):

u just said it backwards

OpenStudy (zaynab123):

x + (x + 4) = 52

OpenStudy (anonymous):

ty I also need help with one more

OpenStudy (xapproachesinfinity):

first name the two number x and y sum is 52 so x+y=52

OpenStudy (xapproachesinfinity):

that's the first step

OpenStudy (xapproachesinfinity):

now we look a the second information

OpenStudy (anonymous):

mhmm

OpenStudy (xapproachesinfinity):

lets take y to be the greater number we know from the problem that the greater number is 4 more than the smaller so we write it y=x+4 then we have x+(x+4)=52

OpenStudy (anonymous):

ok I'm getting it

OpenStudy (xapproachesinfinity):

now we just need to find the smaller number x take off parenthesis x+x+4=52 2x=48 x=24 smaller number 24 bigger number 28

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

finding the actual number is just extra step from me you only need x+(x+4)=52

OpenStudy (anonymous):

ohhh I get it

OpenStudy (anonymous):

can u help with 1 more? it is Natalie works in a toy shop and earns $43 per day. She earns an extra $3 for each toy she sells. If Natalie wants to earn at least $70 per day, which inequality shows the minimum number of toys, n, that she should sell? 43 + 3n ≥ 70, so n ≥ 9 43 + 3n ≤ 70, so n ≤ 9 43 + 3n ≥ 70, so n ≥ 24 43 + 3n ≤ 70, so n ≤ 24

OpenStudy (xapproachesinfinity):

the least means >= ok

OpenStudy (xapproachesinfinity):

she earns 3$ for each toy let n be a number of toys she could sell then we can do 3n to know what she get for each toy sold and we also know that her regular salary is 43$ so 43+3n

OpenStudy (xapproachesinfinity):

we need that to be at least 70$ or up so 43+3n>=70

OpenStudy (xapproachesinfinity):

solve for n

OpenStudy (xapproachesinfinity):

n>=9

OpenStudy (xapproachesinfinity):

you got it!!

OpenStudy (anonymous):

ty, how can I metal you?

OpenStudy (xapproachesinfinity):

hit best response :) you can write testimonial too

OpenStudy (anonymous):

:) have a nice day!

OpenStudy (xapproachesinfinity):

you too:) good luck